OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the points (3, 1) and (2, 9)? y = x +

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

find the slope

Mehek (mehek14):

\(\dfrac{9-1}{2-3}\)

OpenStudy (anonymous):

-8/1

Mehek (mehek14):

so -8

Mehek (mehek14):

now you have y = -8x + ___

Mehek (mehek14):

let's put b in the place after the + y = -8x + b you are going to solve for b let's use the ordered pair (3,1) \(1 = -8*3 + b\)

Mehek (mehek14):

what is \(-8*3\)

OpenStudy (anonymous):

-24

Mehek (mehek14):

\(1 = -24 + b \\add~24\\ 24 + 1 = b\)

OpenStudy (anonymous):

25

Mehek (mehek14):

correct so now your equation is \(y=-8x+25\)